PIERI-TYPE FORMULAS FOR MAXIMAL ISOTROPIC GRASSMANNAINS VIA TRIPLE INTERSECTIONS

FRANK SOTTILE

Abstract. We give an elementary proof of the Pieri-type formula in the cohomology of a Grassmannian of maximal isotropic subspaces of an odd orthogonal or symplectic vector space. This proof proceeds by explicitly computing a triple intersection of Schubert varieties. The decisive step is an exact description of the intersection of two Schubert varieties, from which the multiplicities (which are powers of 2) in the Pieri-type formula are immediately obvious.